by TEI TEI

11/16/07 – TEI received a hard copy of the approved QCE change order in the mail today. TTC filed it for archiving with the project folder. Mike Cavanaugh called to inquire about payment of this invoice and TTC referred him to Bert Watts because all invoices should be paid on this job by this point. TTC is ready to archive this project.

11/06/07 – CRA forwarded Change Order #2 to Bert Watts.

10/30/07 – CRA issued Change Order #2 to delete the allowance balance and forwarded to Kramer & Feldman.

10/26/07 – CRA received final closeout documents.

10/18/07 – MD sent back final change order for signatures on 3 originals. TEI printed and sent 3 new originals to QCE for signing.

10/04/07 – TTC and LKV revised CO4 and sent it to QCE for signing. The new amounts were gleaned from the previous emails and signed change orders.

10/01/07 – TTC received, reviewed and signed payapps #5 and #6 from QCE for this project. I read through all the email correspondences and believe these final two payapps to be correct. Assuming that the change orders that the owner had in hand were signed and correct, the amounts of these pay applications should correspond to requested amounts by the Owner. TEI will mail them to Bert Watts for processing today.

09/28/07 – Awaiting final closeout documents. CRA needs As-built drawings.

09/04/07 – Items mentioned in previous entry were sent to QCE this morning.

08/31/07 – Pay application #05 and #06 need to be updated to show CO#04 instead of CO #05 in the work description (there were only 4 CO’s); the dollar amounts shown are accurate. TEI will correct the dollar amounts shown on CO#04 and send 3 originals to QCE for their signature. TEI will also send 3 original copies of allowance clarification #01 for the entire allowance ($2500) to QCE for their signature. When QCE has received and signed all 6 documents they will send them along with the updated pay application #05 and #06 to TEI. All of the above mention documents will be reviewed, sorted, and sent to HCFD. The documents will be sent to QCE on Tuesday, September 4.

08/23/07 – TEI printed 3 more originals for Change Order #5, signed them, and sent them to QCE with instruction to sign them and forward them to HCFD. TEI also sent the O&M manuals along with the final pay application and closeout documents to HCFD Attn: Bert Watts.

08/17/07 – TEI is awaiting the O&M manuals from QCE in order to close this project out. NR contacted Greg at QCE on August 16, 2007 in order to locate these binders. Greg said he sent them to TEI already but we have no record of receiving them. NR will contact QCE today and have them send another copy.

07/31/07 – TTC approved Change Order and PayApp # 5 for this project but asked for backup data for the change order and requested

07/30/07 – QCE is not responsible for completing the ceiling patch work. The remaining ceiling patch work is going to be completed by the Hamilton County trades department. QCE has completed their portion of work. TEI is awaiting the resubmitted final pay application and close out documents in order to close this project out. TTC emailed the County and QCE for closeout documents on this project and approved their last payapp and told CRA the same info.

07/09/07 – TE still awaiting modified change order from QCE.

06/29/07 – QCE has been informed that they will not be paid for the ceiling and wall patching. TE is awaiting a revised change order and final paperwork from QCE.

06/15/07 – Still awaiting response from QCE.

06/15/07 – TE placed call to Queen City Electric about inaccurate Change Order #04. Awaiting response from QCE.

06/08/07 – Change Order #04 was sent to Hamilton County Facilities Department but had inaccurate information. The change order did not reflect the three previous dollar amounts. The paperwork was given to TA of TE in order to straighten the numbers out.

06/01/07 – TE is still awaiting response from contractors about the punch list. New contract dollar amounts between the contractor and Hamilton County Facilities Department are being clarified. This should be resolved in the next couple of weeks.

05/25/07 – TE issued punch list. Awaiting response from contractors stating that all items have been completed.

05/21/07 – TE completed punch list. Issues with lamps have been corrected. Punch list will be issued early this week.

04/27/07 – Fixtures are 99% (one recessed can yet to be installed). TE has visited site on 2 occasions this week. Project is expected to be complete and TE finish their punch list by the end of next week, May 4 of 2007.

04/20/07 – Project is moving along as expected.

04/13/07 – Light fixtures are currently being installed. Project should be complete by the end of the week of April 20 of 2007.

04/09/07 – Phil stopped down at the site on Thursday and resolved some clarification issues on the type of lights the supplier purchased for this job. TE instructed them to send them back and get the correct cuts for remodel 5″ cans to fit into the drywall soffit.

04/05/07 – QCE has begun installing the fixtures into the space and then we can close this job out.

03/30/07 – Fixtures are scheduled to be installed next Wednesday, April 4 of 2007.

03/23/07 – Project progressing as expected.

03/20/07 – Lights are scheduled to arrive next week and be installed next week by Queen City.

02/19/07 – TEI is waiting for Mindy to get back for funding approval but lighting changes are happening now. Check CRA website for details.

01/10/07 – Const Mtg was held but TEI wasn’t in attendance. Project updates are available on the CRA website.

12/26/06 – TTC moved project over to Mark Donnelly and will get an update from Phil and Spencer this week.

02/06/06 – Spencer submitted a revised floor plan with a new sink location in the back corner of the building. TTC and SBO will pipe up the sink and drain while PEL finishes the electric work. We are shooting to have this job on the street by the end of February 20th for this job.

08/04/05 – Had design meeting today. Discussed schedule and typical workstation layouts, phasing of the project and construction times. C+R will issue official meeting notes for the job. They are talking about moving in around November so design documents have to be collected pretty quickly.

06/24/05 – TTC and Spencer visited the site today to look at the lighting and ceiling heights. TTC brought pictures back for Phil.

06/16/05 – Cost estimates were given to client and we are awaiting direction from the County and C+R before proceeding. TE hasn’t heard anything since 03/21 on moving forward with this project. NO HVAC or Plumbing were included in the cost estimate to the client.

03/01/05 – SOSS Written. TA to approve and send to DES.